1. cement sales were recorded at 1.5 million bags when the price was P500 per bag. When the price decreased to P450 per bag, sales increased to 2 million bags. Cement production on the other hand increased from 1.75 million to 3 million bags when the price increased from P475 to P550 per bag. Based on these figures:

 

  • Estimate the demand and supply equations
  • Determine the market-clearing (equilibrium) price and volume traded.
  • If a price control of P400 is imposed by the government, determine the impact on the market.
  • What kind of price control was imposed by the government.
  • Illustrate graphically.
  • Given the solution to question 1.1, determine the impact on the market if demand increased by 35%.
  • Illustrate graphically.
1
Expert's answer
2021-05-27T11:05:41-0400

(1)sales(demand)

"y=mx+b"

y=price

m=slope

x=quantity

b=y intercept

"slope=\frac{\delta y}{delta x}=\frac{500-450}{2-1.5}=-100"

y intercept

"0=-100(2)+b=200"

demand equation will be

"Q_d=2-\frac{1}{100}p"

production(supply)

"slope=\frac{\delta y}{\delta x}=\frac{550-475}{3-1.75}=60"

y intercept

"0=\u2212180+b=-180"

demand equation will be

"Q=2-\frac{1}{100}p_d"

supply equation will be

"Q_s=3+\frac{1}{60}p"


(2)"Q_d=Q_s"

"2-\frac{1}{100}p=3+\frac{1}{60}p \\-\frac{1}{100}p-\frac{1}{60}=3-2\\\frac{2}{75}=1"


"p=37.5"


"Q=2-\frac{1}{100}p\\=2-\frac{1}{100}(37.5)\\=2-0.375\\=1.625"


(3) The price control will lead to affordable cement though this can reduce the supply of the cement.

This is the price ceiling. the 400 is the highest price that should be charged for cement.



(4)Increase in demand will lead to more quantity of cement demanded which will also lead to a rise in price.